OpenStudy (anonymous):

An electric trimmer blade has rotational inertia 1.55×10−4 kg*m^2. What torque is needed to accelerate this blade from rest to 620 rpm in 1.35 s? How much work was done to accelerate the blade?

OpenStudy (anonymous):

i already found the torque which is 7.45*10^-4 but i dont know how to find the second part which asks how much work was done to accelerate the blade

OpenStudy (experimentx):

F = ma F = torque accn = find angular accn mass = inertia

OpenStudy (anonymous):

Torque is 7.45*10^-4 N*m

OpenStudy (experimentx):

angular acceleration = ??

OpenStudy (anonymous):

it doesn't say.. the question only says "How much work was done to accelerate the blade?" so i don't know if they want the angular acceleration

OpenStudy (experimentx):

you must find it first ... else you can't find the torque

OpenStudy (experimentx):

620 rpm = angular velocity = 620*2*pi/60 = how much is that??

OpenStudy (anonymous):

64.93

OpenStudy (experimentx):

now angular acceleration = (620*2*pi/60-0)/1.35 /s^2, now use this in the relation T = I*a where T is torque, and a is angular accn and I is Inertia

OpenStudy (experimentx):

Jeez .... you surprise me ... finding torque without angular accn

OpenStudy (anonymous):

\[\omega=620*2\pi =64.93 rad/\sec\] \[\alpha=\omega/t= 64.93/1.35= 48.09 rad/\sec^2\] \[\tau=I*\alpha= 1.55*10^-4*48.09= 7.45*10^-3 N.m\] this is how i found torque but i dont know how to find acceleration. what equation am i suppose to use?

OpenStudy (experimentx):

alpha is your angular accn

OpenStudy (experimentx):

that is exactly the equation you are supposed to use

OpenStudy (anonymous):

ok thanks